A small plastic sphere is charged to -10 nC. It is held 1.0 cm above a small glass bead at rest on a table. The bead has a mass of 15 mg and a charge of +10 nC. Will the glass bead “leap up” to the plastic sphere?


1
Expert's answer
2022-06-14T08:42:04-0400

"\frac{F_1}{F_2}=\frac{kq_1q_2}{r^2mg}=60>1," yes.